OpenStudy (anonymous):

Solve –4(x + 2) = 16

OpenStudy (unklerhaukus):

divide by negative four on both side take away two form both sides

OpenStudy (anonymous):

-4(x+2)=16 divide both sides by -4 -4(x+2)/4=-16/4 x+2=-4 subtract 2 from both sides x+2-2=-4-2 x=-6

OpenStudy (jiteshmeghwal9):

distribute -4 btw (x+2) as -4(x)+(-4(2))=16 -4x+(-8)=16 -4x-8=16 -4x=16+8 -4x=24 x=24/-4=-6.